How do police determine how fast a car had been traveling after an accident has occurred? The formula r=25L can be used to approximate the speed r, in miles per hour, of a car that has left a skid mark of length L, in feet. How far will a car skid at 55 mph? at 90 mph?
Answer by malaydassharma(59)

r=25L where speed r, in miles per hour, skid mark of length L in feet
Or, L=r/25
car skid at 55 mph
L=r/25=55/25=2.2 feet
car skid at 90 mph
L=r/25=90/25=3.6 feet